If the force of gravity between 2 identical basketballs (the balls are touching) is 5.12 x 10^-10N, what is the mass of the basketball if the diameter of a basketball is 25 cm.
1
Expert's answer
2020-05-08T16:21:24-0400
"F=G\frac{4m^2}{d^2}\to (5.12 \cdot 10^{-10})=(6.67 \cdot 10^{-11})\frac{4m^2}{0.25^2}"

"m=0.346\ kg"